OpenStudy (anonymous):

Let f be an entire function over the complex plane. Suppose that \[ |f(z)| \le 5 |z|, \,\, \text { for } \, |z|\ge 10 \] What can you say about f?

OpenStudy (anonymous):

ill say one, f is bounded above by 5|z| :p

OpenStudy (anonymous):

You can actually show that f is a very specific function, like a polynomial of a certain degree.

OpenStudy (anonymous):

is there a limit on the degree tho?

OpenStudy (anonymous):

You have to show that it cannot exceed certain degree.

OpenStudy (anonymous):

oh ye, sry misread thought u meant special function, would love to see it but no clue wat to start :P is there a answer somewhere?

OpenStudy (anonymous):

Have you had complex analysis?

OpenStudy (anonymous):

no, well i'll take a look later, ty

OpenStudy (anonymous):

I am sorry for the newbie question, I didn't take any analysis course yet, but can you rewrite f(z) as f(a,b) = 5*(sqrt(a^2 + b^2)) if z = a + bi, and then the inequality becomes: sqrt(a^2 + b^2) >= 10. Don't know if this is the way to prove it, but I was wondering if you are allowed to change a complex-valued function with a real-valued function if you take the modulus of z.

OpenStudy (anonymous):

You need to have a complex analysis course to deal with such a question. Here a hint for people with the above background: Notice that f(0) =0, then f(z) = z g(z) where g(z) is also entire, Show that g(z) is bounded on the whole complex plane and conclude.

OpenStudy (anonymous):

\[f(z).(f(z))^{*} \le 50\] f(z)=x+yi->f(z).f(z)*=x²+y² <=50 so it image lies in a circle with radius sqrt(50)

OpenStudy (anonymous):

This follows from Liouville's Theorem, second corollary on the wiki page: en.wikipedia.org/wiki/Liouville's_theorem_(complex_analysis)

OpenStudy (anonymous):

3rd corollary is related too.